# Graphical models for topological groups: A case study on countable Stone spaces

### v2

@@ Abstract
By analogy with the Cayley graph of a group with respect to a finite generating set or the Cayley–Abels graph of a totally disconnected, locally compact group, we introduce countable connected graphs associated to Polish groups that we term Cayley–Abels–Rosendal graphs. A group admitting a Cayley–Abels–Rosendal graph acts on it continuously, coarsely metrically properly and cocompactly by isometries of the path metric. By an expansion of the Milnor–Schwarz lemma, it follows that the group is generated by a coarsely bounded set and the group equipped with a word metric with respect to a coarsely bounded generating set and the graph are quasi-isometric. In other words, groups admitting Cayley–Abels–Rosendal graphs are topological analogues of finitely generated groups. Our goal is to introduce this topological perspective to a geometric group theorist.
@@ Acknowledgments
Thank you to Christian Rosendal for pointing out an omission in the first version of this paper. Thank you to the referee for helpful comments that have improved the exposition. GD was partially supported by NSF DMS-2303262. BB was partially supported by NSF DMS-1839968. HH was supported by NSF DMS-2303365. RAL was partially supported by NSF DMS-2202942.
@@ 3 Cayley–Abels–Rosendal graphs
A connected, countable simplicial graph $\Gamma$ is a Cayley–Abels–Rosendal graph for a topological group $G$ if $G$ admits a continuous, vertex-transitive action on $\Gamma$ with finitely many orbits of edges and coarsely bounded (necessarily open) vertex stabilizers.
Note: Our definition follows more closely the definition of a Cayley–Abels graph for a totally disconnected locally compact group, rather than that of a Cayley graph. While groups act vertex-transitively on their Cayley graphs, as commonly defined, there is no assumption on the finiteness of the generating set, nor continuity of the action. We also direct the reader to [Ros22, Section 6.2] in which Rosendal carries out a similar discussion in the context of non-Archimedean Polish groups and automorphisms of countable first-order structures.
We begin this section with some generalities on continuous actions of groups on graphs. Suppose that a group $G$ acts on a graph $\Gamma$ continuously. This action induces a representation from $G$ into the group of bijections of the vertex set $V\Gamma$. This latter group is a topological group with the permutation topology, where pointwise stabilizers of finite sets are basic open neighborhoods of the identity, so this representation will be continuous if and only if the stabilizer of a vertex is open in $G$. Such a permutation actually arises as a graph automorphism just when it preserves adjacency.
@@
Now when $\Gamma$ is a simplicial graph and $G$ acts continuously with one orbit of vertices, write $V$ for the stabilizer of the vertex $v$, and notice that there exists an element $k\in G$ sending the oriented edge $e=(g.v,h.v)$ to the oriented edge $e^{\prime}=(g^{\prime}.v,h^{\prime}.v)$ precisely when the pair $(kgV,khV)$ is equal to the pair $(g^{\prime}V,h^{\prime}V)$, or put another way, when the double coset $Vg^{-1}hV$ is equal to the double coset $Vg^{\prime-1}h^{\prime}V$. So when $\Gamma$ is simplicial, every orbit of oriented edges corresponds (uniquely) to a double coset of $V$ in $G$.
Let $A$ be a set of representatives for the double cosets corresponding to oriented edges in $\Gamma$. We may and will assume that $A$ is symmetric, so that $AV=VA$, since $V$ and $A$ are symmetric. Now, notice that if $v$ is the vertex of $\Gamma$ with stabilizer $V$, the vertices $a.v$ are adjacent to $v$. In fact, we have $VA=VAV$. The containment $VA\subset VAV$ is obvious since $V$ is a subgroup. For any element $g\in VAV$, notice that $(v,g.v)$ is an oriented edge of $\Gamma$, since it corresponds to the double coset $VgV$, which is represented by some $a\in A$, since $g\in VAV$. In other words, the set of elements sending $v$ to an adjacent vertex is precisely $VAV$. But it is also $VA$, since the action of $G$ on $V\Gamma$ corresponds to the action of $G$ on $G/V$.
@@
Conversely, given an open subgroup $V\leq G$ and a set of double cosets in $V\backslash G/V$ represented by the set $A$ chosen to be symmetric and meet $V$ trivially, we construct a simplicial graph $\Gamma=\Gamma(V,A)$. Its vertex set is $G/V$ and two vertices (cosets) $gV$ and $hV$ are declared adjacent when the double coset $Vg^{-1}hV$ has a representative in $A$. The action of $G$ on $\Gamma(V,A)$ is continuous, vertex transitive, and $A$ is in one-to-one correspondence with the set of oriented edge orbits.
@@
It is not hard to see that the graph $\Gamma$ will be connected precisely when $AV=VA=VAV$ is an open generating set for $G$. Notice that any continuous vertex-transitive action of a group $G$ on a connected graph $\Gamma$ becomes cobounded as soon as we associate to $\Gamma$ the geodesic path metric that assigns length $1$ to each edge. If such an action is to be coarsely metrically proper, it must be the case that the set $VAV=AV=VA$ associated to $\Gamma$ as above is coarsely bounded in $G$, since it fails to move the closed $1$-neighborhood of $v$ off of itself. But since $\Gamma$ is connected, the coarsely bounded open set $VAV$ generates $G$.
@@
Suppose that $G$ admits a Cayley–Abels–Rosendal graph. Then $G$ is boundedly generated and by the Milnor–Schwarz Lemma is quasi-isometric to any Cayley–Abels–Rosendal graph for $G$. If $G$ is boundedly generated and satisfies the hypotheses of 1, then for any open, coarsely bounded subgroup $V$ of $G$ (supposing it exists), there exists a Cayley–Abels–Rosendal graph for $G$ with $V$ as the stabilizer of a vertex.
@@
The foregoing argument shows that if $G$ has a Cayley–Abels–Rosendal graph $\Gamma$, then the graph $\Gamma$ is $G$-equivariantly isomorphic to $\Gamma(V,A)$ for some open coarsely bounded subgroup $V$ and set of double coset representatives $A$. The set $VAV$ generates $G$ since $\Gamma(V,A)$ is connected, as does $V\cup A$, which is coarsely bounded since there are finitely many orbits of edges; hence $A$ is a finite set.
Supposing that $G$ is boundedly generated and satisfies the hypotheses of 1 and that $V$ is an open, coarsely bounded subgroup, we may apply Rosendal’s Criterion to a coarsely bounded generating set $S$ of $G$ with $V$ as our open identity neighborhood. This provides a finite set $F$ and $N\in\mathbb{N}$ such that $S\subset(FV)^{N}$. We may without loss assume that $F$ is symmetric and is disjoint from $V$. Since $S\subset(FV)^{N}$, we have that $VFV$ generates $G$. Therefore, the graph $\Gamma=\Gamma(V,F)$ is connected and admits a continuous, vertex-transitive action of $G$ with finitely many orbits of edges (the finite set $F$ is in one-to-one correspondence with the set of oriented edge orbits) and vertex stabilizers conjugate to the coarsely bounded subgroup $V$. $\Gamma$ is thus a Cayley–Abels–Rosendal graph for $G$. ∎
@@ 4.2 Coarsely Bounded and Locally Bounded
Suppose $X$ is a self-similar Stone space, and that $U$ is an open subset of $\operatorname{Homeo}(X)$. Let $\mathcal{P}=P_{0}\sqcup P_{1}\sqcup\cdots\sqcup P_{n}$ be a clopen partition of $X$ such that $U$ contains $V=\operatorname{Stab}(\mathcal{P})$ and such that $P_{0}$ contains a clopen subset homeomorphic to $X$. There exists an involution $f\notin V$ such that the graph $\Gamma(V,A)$ as defined in Section 3 associated to $V$ and to $A=\{f\}$ has diameter at most three (so is connected) and admits a vertex-transitive and edge-transitive action of $\operatorname{Homeo}(X)$.
@@
We first construct a graph $\Gamma$ associated to the given partition $\mathcal{P}$ that admits a vertex-transitive and edge-transitive action of $\operatorname{Homeo}(X)$. We show that the graph $\Gamma$ has diameter at most three, and then realize it as combinatorially identical to the graph $\Gamma(V,A)$ as defined in Section 3 associated to the subgroup $V=\operatorname{Stab}(\mathcal{P})$ and the set $A=\{f\}$ for some involution $f\notin V$.
@@
Finally, we see that $\Gamma$ may be identified with the graph $\Gamma(V,A)$ as defined in Section 3 associated to the subgroup $V=\operatorname{Stab}(\mathcal{P})$ and the set $A=\{f\}$ for some involution $f\notin V$. Indeed, a homeomorphism of $X$ fixes the vertex associated to $\mathcal{P}$ precisely when it is contained in $V$, and we saw that any two adjacent vertices are exchanged by some involution and that $\Gamma$ has one $\operatorname{Homeo}(X)$-orbit of edges, so choosing for $f$ any such involution not contained in $V$ (for example, one which exchanges $\mathcal{P}$ itself with an adjacent partition) proves the claim. ∎
@@
We first prove that every orbit in every continuous action of $\operatorname{Homeo}(X)$ by isometries on a metric space is Lipschitz-dominated by an action on a connected graph as in the statement of 13. Suppose that $\operatorname{Homeo}(X)$ acts continuously and by isometries on a metric space $X$. Take for $U$ in the statement above the open set comprising those elements failing to move a point $x\in X$ a distance more than $\epsilon>0$, and construct the graph $\Gamma(V,A)$ with vertex stabilizer associated to $V$ and $A$. Choosing a transversal of coset representatives $\mathcal{T}$ of $V$ defines a map $\Phi\colon V\Gamma\to X$ as $gV\mapsto g.x$ for $g\in\mathcal{T}$. Write $\delta$ for $d(x,f.x)$, where $f$ is the involution in the statement of the proposition. The map $\Phi$ is Lipschitz: Let $gV$ and $hV$ be connected by an edge in $\Gamma(V,A)$ so that $Vg^{-1}hV=VAV$. Then
@@
Since every such graph $\Gamma(V,A)$ has diameter at most three, we conclude. ∎
@@ 4.3 Cayley-Abels-Rosendal Graphs for Successor Ordinals
Now, recall that for any good partition $\mathcal{P}$, the stabilizer $\operatorname{Stab}(\mathcal{P})$ is open and coarsely bounded. The group $\operatorname{Homeo}(X_{\alpha,n})$ acts continuously and transitively on the set of good partitions. Fix a good partition $\mathcal{P}$ and a maximal shift $g\in\operatorname{Homeo}(X_{\alpha,n})$ which moves $\mathcal{P}$, and let $A=\{g^{\pm 1}\}$.
@@
The graph $\Gamma$ is of the form $\Gamma(\operatorname{Stab}(\mathcal{P}),A)$, where $A=\{g^{\pm 1}\}$ and is thus a Cayley–Abels–Rosendal graph for $\operatorname{Homeo}(X_{\alpha,n})$.
@@
By the paragraph above, the group $\operatorname{Homeo}(X_{\alpha,n})$ acts transitively on $\Gamma(\alpha,n)$ with stabilizers conjugate to $\operatorname{Stab}(\mathcal{P})$, which is open in $\operatorname{Homeo}(X_{\alpha,n})$ and coarsely bounded. Now, supposing that $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are good partitions of $X$ which differ from $\mathcal{P}$ by a shift, notice that there is a homeomorphism preserving $\mathcal{P}$ taking one of the shifted sets, which is homeomorphic to $X_{\beta,1}$, to the other. See Figure 6 for an example of such a homeomorphism. Thus the group $\operatorname{Homeo}(X_{\alpha,n})$ acts edge-transitively on the graph $\Gamma$, and $A$ is therefore a complete set of representatives for the orbits of oriented edges of $\Gamma$.

### v3

@@ Abstract
By analogy with the Cayley graph of a group with respect to a finite generating set or the Cayley–Abels graph of a totally disconnected, locally compact group, we detail countable connected graphs associated to Polish groups that we term Cayley–Abels–Rosendal graphs. A group admitting a Cayley–Abels–Rosendal graph acts on it continuously, coarsely metrically properly and cocompactly by isometries of the path metric. By an expansion of the Milnor–Schwarz lemma, it follows that the group is generated by a coarsely bounded set and the group equipped with a word metric with respect to a coarsely bounded generating set and the graph are quasi-isometric. In other words, groups admitting Cayley–Abels–Rosendal graphs are topological analogues of finitely generated groups. Our goal is to introduce this topological perspective on the work of Rosendal to a geometric group theorist.
@@ 1 Introduction
Let $X_{\alpha,n}$ be a countable Stone space with $\alpha>0$. The group $\operatorname{Homeo}(X_{\alpha,n})$ is always locally bounded and is
@@
When $\alpha=0$, $X_{\alpha,n}$ is a finite discrete set, so its homeomorphism group is coarsely bounded.
@@ Acknowledgments
Thank you to Christian Rosendal and Carolyn Abbott for pointing out two omissions in the first version of this paper. Thank you to the referee for helpful comments that have improved the exposition. GD was partially supported by NSF DMS-2303262. BB was partially supported by NSF DMS-1839968. HH was supported by NSF DMS-2303365. RAL was partially supported by NSF DMS-2202942.
@@ 3 Cayley–Abels–Rosendal graphs
The results in this section are inspired by [Ros22, Section 6.2] and [Led22, Section 2]. There the authors work in the settings of non-Archimedean Polish groups (the closed subgroups of the symmetric group on a countable set) or totally disconnected, locally compact groups. In each case, the groups considered have a wealth of open subgroups, making them well-suited to the constructions we describe here. Our perspective is broadly very similar. One difference is that since we work in greater generality, our results show that their constructions are in some sense the only ones available.
First, a definition. A connected, countable simplicial graph $\Gamma$ is a Cayley–Abels–Rosendal graph for a topological group $G$ if $G$ admits a continuous, vertex-transitive action on $\Gamma$ with finitely many orbits of edges and coarsely bounded (necessarily open) vertex stabilizers.
Note: Our definition follows more closely the definition of a Cayley–Abels graph for a totally disconnected locally compact group, rather than that of a Cayley graph. While groups act vertex-transitively on their Cayley graphs, as commonly defined, there is no assumption on the finiteness of the generating set, nor continuity of the action.
Next we turn to some generalities on continuous actions of groups on graphs. Suppose that a group $G$ acts on a graph $\Gamma$ continuously. This action induces a representation from $G$ into the group of bijections of the vertex set $V\Gamma$. This latter group is a topological group with the permutation topology, where pointwise stabilizers of finite sets are basic open neighborhoods of the identity, so this representation will be continuous if and only if the stabilizer of a vertex is open in $G$. Such a permutation actually arises as a graph automorphism just when it preserves adjacency.
@@
Now when $\Gamma$ is a simplicial graph and $G$ acts continuously with one orbit of vertices, write $V$ for the stabilizer of a fixed vertex $v$, and notice that there exists an element $k\in G$ sending the oriented edge $e=(g.v,h.v)$ to the oriented edge $e^{\prime}=(g^{\prime}.v,h^{\prime}.v)$ precisely when the pair $(kgV,khV)$ is equal to the pair $(g^{\prime}V,h^{\prime}V)$, or put another way, when the double coset $Vg^{-1}hV$ is equal to the double coset $Vg^{\prime-1}h^{\prime}V$. So when $\Gamma$ is simplicial, every orbit of oriented edges corresponds (uniquely) to a double coset of $V$ in $G$.
Let $U$ be the (open) set of elements $g$ in $G$ such that $g.v$ is either equal to $v$ or adjacent to it. Observe that $VUV=U$: if $g\in U$ and $h$, $k\in V$, first we note that we have $hgk.v=hg.v$. If $g.v=v$, then $hg.v=v$, so $hgk\in U$. On the other hand, if $g.v=w$ which is adjacent to $v$, then since $h.v=v$, the element $h$ must send $w$ to another vertex adjacent to $v$, whence $hgk\in U$.
Thus if $g\in U$, then actually its whole coset $gV$ is contained in $U$. Let $A$ be a set of representatives for the cosets in $G/V$ contained in $U$. Since $U$ is symmetric by definition, we may choose $A$ to be symmetric, whence $U=VUV=AV={(VA)}^{-1}=V^{-1}A^{-1}=VA$, since $V$ is a subgroup. When $G$ is alternately Polish or $G$ is locally compact and $U$ is compact, the set $A$ is correspondingly countable or finite, and is in one-to-one correspondence with the set of vertices of $\Gamma$ equal to or adjacent to $v$. Thus when $\Gamma$ is simplicial, $A-\{1\}$ is in one-to-one correspondence with the set of edges incident to $v$.
Let $F$ be a set of representatives for the double cosets corresponding to oriented edge orbits in $\Gamma$. We choose the notation $F$ to follow Rosendal [Ros22] and note that in the case that $\Gamma$ is a Cayley–Abels–Rosendal graph, $F$ will be a finite set. We may choose $F$ such that $F\subset(A-\{1\})$ and such that $F$ is symmetric.
@@
Conversely, given an open subgroup $V\leq G$ and a set of double cosets in $V\backslash G/V$ represented by the set $F$ chosen to be symmetric and meet $V$ trivially, we construct a simplicial graph $\Gamma=\Gamma(V,F)$. Its vertex set is $G/V$ and two vertices (cosets) $gV$ and $hV$ are declared adjacent when the double coset $Vg^{-1}hV$ has a representative in $F$. Put more plainly, we declare that the vertex (coset) $gV$ is adjacent to $gafV$ for each $a\in V$ and $f\in F$. The set of edges of the simplicial graph $\Gamma$ incident to the vertex $V$ is thus in bijection with the set of cosets $gV$ contained in the open set $VFV$ (notice that this set is likely much smaller than the set of elements $VF$). The action of $G$ on $\Gamma(V,F)$ is continuous, vertex transitive, and $F$ is in one-to-one correspondence with the set of oriented edge orbits.
@@
It is not hard to see that the graph $\Gamma$ will be connected precisely when $VFV\cup V$ (which is equal to $U=AV=VA$ in the discussion above) is an open generating set for $G$. Notice that any continuous vertex-transitive action of a group $G$ on a connected graph $\Gamma$ becomes cobounded as soon as we associate to $\Gamma$ the geodesic path metric that assigns length $1$ to each edge. If such an action is to be coarsely metrically proper, it must be the case that the set $VFV\cup V$ associated to $\Gamma$ as above is coarsely bounded in $G$, since it fails to move the closed $1$-neighborhood of $v$ off of itself. But since $\Gamma$ is connected, the coarsely bounded open set $VFV\cup V$ generates $G$.
@@
Suppose that $G$ admits a Cayley–Abels–Rosendal graph. Then $G$ is boundedly generated and by the Milnor–Schwarz Lemma is quasi-isometric to any Cayley–Abels–Rosendal graph for $G$. If $G$ is boundedly generated and satisfies the hypotheses of 1, then for any open, coarsely bounded subgroup $V$ of $G$ (supposing one exists), there exists a Cayley–Abels–Rosendal graph for $G$ with $V$ as the stabilizer of a vertex.
@@
The foregoing argument shows that if $G$ has a Cayley–Abels–Rosendal graph $\Gamma$, then the graph $\Gamma$ is $G$-equivariantly isomorphic to $\Gamma(V,F)$ for some open coarsely bounded subgroup $V$ and finite set of double coset representatives $F$. The set $VFV$ generates $G$ since $\Gamma(V,F)$ is connected, as does $V\cup F$, which is coarsely bounded since $F$ is finite.
Supposing that $G$ is boundedly generated and satisfies the hypotheses of 1 and that $V$ is an open, coarsely bounded subgroup, we may apply Rosendal’s Criterion to a coarsely bounded generating set $S$ of $G$ with $V$ as our open identity neighborhood. This provides a finite set $F$ and $N\in\mathbb{N}$ such that $S\subset(FV)^{N}$. We may assume that $F$ is symmetric and is disjoint from $V$. Since $S\subset(FV)^{N}$, we have that $VFV$ generates $G$. Therefore, the graph $\Gamma=\Gamma(V,F)$ is connected and admits a continuous, vertex-transitive action of $G$ with finitely many orbits of edges (the finite set $F$ is in one-to-one correspondence with the set of oriented edge orbits) and vertex stabilizers conjugate to the coarsely bounded subgroup $V$. The graph $\Gamma$ is thus a Cayley–Abels–Rosendal graph for $G$. ∎
@@ 4.2 Coarsely Bounded and Locally Bounded
Suppose $X$ is a self-similar Stone space, and that $U$ is an open subset of $\operatorname{Homeo}(X)$. Let $\mathcal{P}=P_{0}\sqcup P_{1}\sqcup\cdots\sqcup P_{n}$ be a clopen partition of $X$ such that $U$ contains $V=\operatorname{Stab}(\mathcal{P})$ and such that $P_{0}$ contains a clopen subset homeomorphic to $X$. There exists an involution $f\notin V$ such that the graph $\Gamma(V,F)$ as defined in Section 3 associated to $V$ and to $F=\{f\}$ has diameter at most three (so is connected) and admits a vertex-transitive and edge-transitive action of $\operatorname{Homeo}(X)$.
@@
We first construct a graph $\Gamma$ associated to the given partition $\mathcal{P}$ that admits a vertex-transitive and edge-transitive action of $\operatorname{Homeo}(X)$. We show that the graph $\Gamma$ has diameter at most three, and then realize it as combinatorially identical to the graph $\Gamma(V,F)$ as defined in Section 3 associated to the subgroup $V=\operatorname{Stab}(\mathcal{P})$ and the set $F=\{f\}$ for some involution $f\notin V$.
@@
Finally, we see that $\Gamma$ may be identified with the graph $\Gamma(V,F)$ as defined in Section 3 associated to the subgroup $V=\operatorname{Stab}(\mathcal{P})$ and the set $F=\{f\}$ for some involution $f\notin V$. Indeed, a homeomorphism of $X$ fixes the vertex associated to $\mathcal{P}$ precisely when it is contained in $V$, and we saw that any two adjacent vertices are exchanged by some involution and that $\Gamma$ has one $\operatorname{Homeo}(X)$-orbit of edges, so choosing for $f$ any such involution not contained in $V$ (for example, one which exchanges $\mathcal{P}$ itself with an adjacent partition) proves the claim. ∎
@@
We first prove that every orbit in every continuous action of $\operatorname{Homeo}(X)$ by isometries on a metric space is Lipschitz-dominated by an action on a connected graph as in the statement of 13. Suppose that $\operatorname{Homeo}(X)$ acts continuously and by isometries on a metric space $X$. Take for $U$ in the statement above the open set comprising those elements failing to move a point $x\in X$ a distance more than $\epsilon>0$, and construct the graph $\Gamma(V,F)$ with vertex stabilizer associated to $V$ and $F$. Choosing a transversal of coset representatives $\mathcal{T}$ of $V$ defines a map $\Phi\colon V\Gamma\to X$ as $gV\mapsto g.x$ for $g\in\mathcal{T}$. Write $\delta$ for $d(x,f.x)$, where $f$ is the involution in the statement of the proposition. The map $\Phi$ is Lipschitz: Let $gV$ and $hV$ be connected by an edge in $\Gamma(V,F)$ so that $Vg^{-1}hV=VfV=VFV$. Then
@@
Since every such graph $\Gamma(V,F)$ has diameter at most three, we conclude. ∎
@@ 4.3 Cayley-Abels-Rosendal Graphs for Successor Ordinals
Now, recall that for any good partition $\mathcal{P}$, the stabilizer $\operatorname{Stab}(\mathcal{P})$ is open and coarsely bounded. The group $\operatorname{Homeo}(X_{\alpha,n})$ acts continuously and transitively on the set of good partitions. Fix a good partition $\mathcal{P}$ and a maximal shift $g\in\operatorname{Homeo}(X_{\alpha,n})$ which moves $\mathcal{P}$, and let $F=\{g^{\pm 1}\}$.
@@
The graph $\Gamma$ is of the form $\Gamma(\operatorname{Stab}(\mathcal{P}),F)$, where $F=\{g^{\pm 1}\}$ and is thus a Cayley–Abels–Rosendal graph for $\operatorname{Homeo}(X_{\alpha,n})$.
@@
By the paragraph above, the group $\operatorname{Homeo}(X_{\alpha,n})$ acts transitively on $\Gamma(\alpha,n)$ with stabilizers conjugate to $\operatorname{Stab}(\mathcal{P})$, which is open in $\operatorname{Homeo}(X_{\alpha,n})$ and coarsely bounded. Now, supposing that $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are good partitions of $X$ which differ from $\mathcal{P}$ by a shift, notice that there is a homeomorphism preserving $\mathcal{P}$ taking one of the shifted sets, which is homeomorphic to $X_{\beta,1}$, to the other. See Figure 6 for an example of such a homeomorphism. Thus the group $\operatorname{Homeo}(X_{\alpha,n})$ acts edge-transitively on the graph $\Gamma$, and $F$ is therefore a complete set of representatives for the orbits of oriented edges of $\Gamma$.